# Tallahassee, FL 32306-4350, USA

### v1

@@ 10 TeV 3 TeV 1 TeV - noRGE
$C_{qe}^{[aabb]}$ $-0.23\pm 0.36$ $10.4$ F $-0.015\pm 0.032$ $10.8$ F $-0.0021\pm 0.0035$ $10.3$ F
$C_{qu}^{(1)[3333]}$ $-3.7\pm 5.8$ $2.6$ F/EW $-0.38\pm 0.80$ $2.2$ F $4.77\pm 2.53$ 0.32
$C_{qu}^{(1)[aa33]}$ ${\color[rgb]{1,0,0}8.9\pm 5.4}$ F $0.95\pm 0.74$ $1.9$ F/EW
$C_{qu}^{(1)[33aa]}$ ${\color[rgb]{1,0,0}-4.9\pm 8.8}$ F $-0.5\pm 1.7$ $1.6$ F $-0.85\pm 2.15$ 0.44 T
$C_{qu}^{(1)[aabb]}$ $-2.7\pm 8.9$ F/DY $0.2\pm 1.5$ $1.7$ DY
$C_{qu}^{(8)[3333]}$ ${\color[rgb]{1,0,0}-10.3\pm 4.7}$ F $-6.2\pm 2.5$ $0.9$ F ${\color[rgb]{1,0,0}-7.6\pm 2.5}$ T
@@
$C_{quqd}^{(1)[3333]}$ $-0.02\pm 0.77$ $7.8$ F $-0.024\pm 0.087$ $6.5$ F
$C_{quqd}^{(8)[3333]}$ $-0.27\pm 0.69$ $7.7$ F $-0.033\pm 0.081$ $6.8$ F
$C_{lequ}^{(1)[3333]}$ $-1.2\pm 3.7$ $3.5$ H $-0.10\pm 0.40$ $3.1$ H
@@
$C$ $\Lambda/\sqrt{|C|}$ $C$ $\Lambda/\sqrt{|C|}$ $C$ $\Lambda/\sqrt{|C|}$ $C$ $\Lambda/\sqrt{|C|}$
$C_{\phi q}^{(1)[33]}$ $0.076\pm 0.081$ $6.1$ $0.07\pm 0.15$ $5.0$ $0.07\pm 0.15$ $5.1$ $0.07\pm 0.15$ $5.0$
$C_{\phi q}^{(1)[aa]}$ $-0.06\pm 0.10$ $5.9$ $0.46\pm 0.59$ $2.3$ $0.53\pm 0.59$ $2.3$ $0.42\pm 0.61$ $2.3$
$C_{\phi q}^{(3)[33]}$ $0.083\pm 0.083$ $6.0$ $0.22\pm 0.17$ $4.0$ $0.24\pm 0.17$ $4.0$ $0.19\pm 0.17$ $4.0$
$C_{\phi q}^{(3)[aa]}$ $0.078\pm 0.055$ $7.0$ $0.123\pm 0.067$ $5.9$ $0.132\pm 0.068$ $5.9$ $0.128\pm 0.068$ $5.9$
$C_{dB}^{[33]}$ $-0.00020\pm 0.00052$ $85.3$ $3.7\pm 5.1$ $0.8$ $0.48\pm 2.14$ $1.3$ $4.4\pm 5.1$ $0.8$
$C_{dW}^{[33]}$ $0.00035\pm 0.00099$ $61.5$ $0.44\pm 0.57$ $2.4$ $-0.005\pm 0.013$ $16.7$ $0.42\pm 0.56$ $2.4$
$C_{dG}^{[33]}$ $-0.004\pm 0.012$ $17.5$ $-2.3\pm 4.0$ $0.9$ $0.12\pm 0.32$ $3.3$ $-2.3\pm 3.9$ $1.0$
$C_{qq}^{(1)[3333]}$ $-0.043\pm 0.020$ $10.7$ $-0.65\pm 0.74$ $2.0$ $-0.40\pm 0.70$ $2.2$ $-0.66\pm 0.76$ $2.1$
$C_{qq}^{(1)[aa33]}$ $0.0192\pm 0.0096$ $15.4$ $-3.1\pm 1.6$ $1.2$ $-2.2\pm 1.1$ $1.4$ $-3.1\pm 1.5$ $1.2$
@@
$C_{qq}^{(1)[aabb]}$ $-0.036\pm 0.018$ $11.2$ $0.7\pm 1.2$ $1.7$ $1.0\pm 1.1$ $1.7$ $0.6\pm 1.2$ $1.7$
$C_{qq}^{(1)[abba]}$ $-0.036\pm 0.021$ $10.1$ $0.24\pm 0.59$ $2.6$ $0.38\pm 0.58$ $2.5$ $0.23\pm 0.59$ $2.5$
$C_{qq}^{(3)[3333]}$ $-0.040\pm 0.020$ $410$±1.4plus-or-minus1.01.41.0$\pm$ 1.41.0 ± 1.4 $1.5$ $1.4\pm 1.2$ $1.5$ $0.6\pm 1.4$ $1.5$
@@
$C_{qq}^{(3)[aabb]}$ $-0.032\pm 0.019$ $11.2$ $0.032\pm 0.078$ $6.9$ $0.031\pm 0.079$ $7.0$ $0.031\pm 0.079$ $7.0$
$C_{qq}^{(3)[abba]}$ $-0.036\pm 0.018$ $11.1$ $0.06\pm 0.20$ $4.5$ $0.06\pm 0.20$ $4.4$ $0.06\pm 0.20$ $4.4$
$C_{lq}^{(1)[aa33]}$ $0.11\pm 0.15$ $4.6$ $0.30\pm 0.32$ $3.1$ $0.30\pm 0.33$ $3.1$ $0.33\pm 0.33$ $3.1$
$C_{lq}^{(3)[aa33]}$ $0.07\pm 0.13$ $5.1$ $-0.02\pm 0.25$ $4.2$ $-0.03\pm 0.25$ $4.3$ $-0.01\pm 0.25$ $4.2$
$C_{qe}^{[33aa]}$ $0.020\pm 0.035$ $9.9$ $-0.58\pm 0.45$ $2.5$ $-0.59\pm 0.40$ $2.6$ $-0.55\pm 0.46$ $2.5$
$C_{qe}^{[aabb]}$ $-0.015\pm 0.032$ $10.8$ $0.055\pm 0.099$ $6.0$ $0.06\pm 0.10$ $6.0$ $0.05\pm 0.10$ $6.0$
$C_{qu}^{(1)[3333]}$ $-0.38\pm 0.80$ $2.2$ $0.6\pm 1.3$ $1.7$ $0.7\pm 1.3$ $1.7$ $0.7\pm 1.3$ $1.7$
$C_{qu}^{(1)[aa33]}$ $0.95\pm 0.74$ $1.9$ $0.9\pm 2.3$ $1.3$ $1.5\pm 2.2$ $1.2$ $1.1\pm 2.2$ $1.2$
$C_{qu}^{(1)[33aa]}$ $-0.5\pm 1.7$ $1.6$ $-2.4\pm 2.2$ $1.1$ $-2.4\pm 2.3$ $1.2$ $-2.5\pm 2.3$ $1.1$
$C_{qu}^{(1)[aabb]}$ $0.2\pm 1.5$ $1.7$ $1.5\pm 1.9$ $1.3$ $1.5\pm 1.8$ $1.3$ $1.3\pm 1.9$ $1.3$
@@ 10 TeV Individual 10 TeV Global
$C$ $\Lambda/\sqrt{|C|}$ $C$ $\Lambda/\sqrt{|C|}$
@@
$C_{\phi u}^{[aa]}$ $4.8\pm 4.3$ $2.8$ ${\color[rgb]{1,0,0}-6.3\pm 8.8}$
$C_{\phi d}^{[33]}$ ${\color[rgb]{1,0,0}-10.9\pm 4.2}$ ${\color[rgb]{1,0,0}-9.8\pm 5.3}$
@@
$C_{ll}^{[aabb]}$ $-2.5\pm 2.6$ $3.7$
$C_{ll}^{[abba]}$ $-0.1\pm 1.0$ $6.8$ $-3.0\pm 3.5$ $3.6$
$C_{qq}^{(1)[3333]}$ $-0.45\pm 0.23$ $10.7$
$C_{qq}^{(1)[aa33]}$ $0.211\pm 0.099$ $16.1$
$C_{qq}^{(1)[a33a]}$ $0.24\pm 0.13$ $14.4$
$C_{qq}^{(1)[aabb]}$ $-0.43\pm 0.22$ $10.2$
$C_{qq}^{(1)[abba]}$ $-0.47\pm 0.22$ $10.7$
$C_{qq}^{(3)[3333]}$ $-0.56\pm 0.23$ $10.2$
$C_{qq}^{(3)[aa33]}$ $0.22\pm 0.13$ $13.5$
$C_{qq}^{(3)[a33a]}$ $0.18\pm 0.11$ $15.7$
$C_{qq}^{(3)[aabb]}$ $-0.36\pm 0.21$ $11.5$
$C_{qq}^{(3)[abba]}$ $-0.40\pm 0.21$ $10.6$
@@
$C_{lq}^{(1)[aa33]}$ $1.4\pm 1.6$ $4.6$ ${\color[rgb]{1,0,0}-7.0\pm 8.1}$
$C_{lq}^{(1)[33aa]}$ $3.6\pm 2.5$ $3.4$ $-2.6\pm 8.2$
$C_{lq}^{(1)[aabb]}$ $-0.08\pm 0.53$ $9.5$ $-3.9\pm 4.6$ $2.8$
$C_{lq}^{(3)[3333]}$ $8.2\pm 5.8$ $2.6$
$C_{lq}^{(3)[aa33]}$ $0.5\pm 1.4$ $5.5$ $5.6\pm 7.0$
$C_{lq}^{(3)[33aa]}$ $-0.21\pm 0.29$ $11.1$ $-0.26\pm 1.1$ $7.5$
$C_{lq}^{(3)[aabb]}$ $-0.017\pm 0.059$ $27.5$ $0.06\pm 0.85$ $7.9$
$C_{ee}^{[aa33]}$ $0.0\pm 2.1$ $4.9$ $-3.7\pm 7.5$ $2.7$
$C_{ee}^{[aabb]}$ $-1.4\pm 1.5$ $4.8$ $-3.7\pm 6.5$ $2.7$
@@
$C_{le}^{[aa33]}$ $-4.2\pm 7.1$ ${\color[rgb]{1,0,0}-7.4\pm 6.0}$
$C_{le}^{[33aa]}$ ${\color[rgb]{1,0,0}-6.7\pm 7.7}$ ${\color[rgb]{1,0,0}-8.4\pm 6.7}$
$C_{le}^{[aabb]}$ $0.4\pm 4.2$ $3.3$ $0.35\pm 5.4$ $3.4$
@@
$C_{qe}^{[33aa]}$ $0.22\pm 0.38$ $10.0$ ${\color[rgb]{1,0,0}8.0\pm 7.0}$
$C_{qe}^{[aabb]}$ $-0.23\pm 0.36$ $10.4$ ${\color[rgb]{1,0,0}-6.1\pm 7.2}$
$C_{qu}^{(1)[3333]}$ $-3.7\pm 5.8$ $2.6$ ${\color[rgb]{1,0,0}7.5\pm 7.5}$
$C_{qu}^{(1)[aa33]}$ ${\color[rgb]{1,0,0}8.9\pm 5.4}$
$C_{qu}^{(1)[33aa]}$ ${\color[rgb]{1,0,0}-4.9\pm 8.8}$
$C_{qu}^{(1)[aabb]}$ $-2.7\pm 8.9$
@@
$C_{quqd}^{(1)[3333]}$ $-0.02\pm 0.77$ $7.8$
$C_{quqd}^{(8)[3333]}$ $-0.27\pm 0.69$ $7.7$
$C_{lequ}^{(1)[3333]}$ $-1.2\pm 3.7$ $3.5$

### v2

@@ 10 TeV 3 TeV 1 TeV - noRGE
$C_{qe}^{[aabb]}$ $0.23\pm 0.36$ $10.4$ F $0.015\pm 0.032$ $10.8$ F $0.0021\pm 0.0035$ $10.3$ F
$C_{qu}^{(1)[3333]}$ $0.3\pm 3.1$ $3.9$ F $0.08\pm 0.36$ $3.4$ F $4.77\pm 2.53$ $0.32$ T
$C_{qu}^{(1)[aa33]}$ $0.1\pm 2.9$ $4.2$ F $0.02\pm 0.34$ $3.6$ F $1.1\pm 3.3$ $0.4$ T
$C_{qu}^{(1)[33aa]}$ $\color[rgb]{1,0,0}-7.0\pm 6.9$} F $2.5\pm 2.3$ $1.1$ F
$C_{qu}^{(1)[aabb]}$ $3.6\pm 9.1$ DY $1.5\pm 1.9$ $1.3$ DY
$C_{qu}^{(8)[3333]}$ $\color[rgb]{1,0,0}-10.3\pm 4.7$} F $6.2\pm 2.5$ $0.9$ F $\color[rgb]{1,0,0}-7.6\pm 2.5$} T
@@
$C_{quqd}^{(1)[3333]}$ $0.10\pm 0.40$ $10.2$ F $0.014\pm 0.047$ $9.2$ F
$C_{quqd}^{(8)[3333]}$ $0.16\pm 0.48$ $9.2$ F $0.020\pm 0.054$ $8.1$ F
$C_{lequ}^{(1)[3333]}$ $1.2\pm 3.7$ $3.5$ H $0.10\pm 0.40$ $3.1$ H
@@
$C_{i}$ $\Lambda/\sqrt{|C_{i}|}$ $C_{i}$ $\Lambda/\sqrt{|C_{i}|}$ $C_{i}$ $\Lambda/\sqrt{|C_{i}|}$ $C_{i}$ $\Lambda/\sqrt{|C_{i}|}$
$C_{\phi q}^{(1)[33]}$ $0.029\pm 0.036$ $9.5$ $0.08\pm 0.14$ $5.0$ $0.06\pm 0.14$ $5.1$ $0.08\pm 0.14$ $5.0$
$C_{\phi q}^{(1)[aa]}$ $0.022\pm 0.037$ $9.5$ $0.46\pm 0.60$ $2.4$ $0.48\pm 0.56$ $2.4$ $0.38\pm 0.60$ $2.4$
$C_{\phi q}^{(3)[33]}$ $0.026\pm 0.030$ $10.1$ $0.21\pm 0.18$ $4.0$ $0.22\pm 0.17$ $4.0$ $0.20\pm 0.18$ $4.0$
$C_{\phi q}^{(3)[aa]}$ $0.010\pm 0.030$ $11.1$ $0.124\pm 0.066$ $6.0$ $0.134\pm 0.068$ $5.8$ $0.122\pm 0.069$ $5.9$
$C_{dB}^{[33]}$ $0.00012\pm 0.00037$ $101.0$ $3.7\pm 5.1$ $0.8$ $\pm 1.7 1.31.3$ $4.4\pm 5.1$ $0.8$
$C_{dW}^{[33]}$ $0.00028\pm 0.00070$ $72.3$ $0.44\pm 0.57$ $2.4$ $0.0033\pm 0.0091$ $20.2$ $0.42\pm 0.56$ $2.4$
$C_{dG}^{[33]}$ $0.0014\pm 0.0031$ $33.6$ $2.3\pm 4.0$ $0.9$ $0.08\pm 0.22$ $4.2$ $2.3\pm 3.9$ $1.0$
$C_{qq}^{(1)[3333]}$ $0.043\pm 0.020$ $10.7$ $0.65\pm 0.74$ $2.0$ $0.40\pm 0.70$ $2.2$ $0.66\pm 0.76$ $2.1$
$C_{qq}^{(1)[aa33]}$ $0.0192\pm 0.0096$ $15.4$ $3.1\pm 1.6$ $1.2$ $2.2\pm 1.1$ $1.4$ $3.1\pm 1.5$ $1.2$
@@
$C_{qq}^{(1)[aabb]}$ $0.036\pm 0.018$ $11.2$ $0.7\pm 1.2$ $1.7$ $1.0\pm 1.1$ $1.7$ $0.6\pm 1.2$ $1.7$
$C_{qq}^{(1)[abba]}$ $0.036\pm 0.021$ $10.1$ $0.24\pm 0.59$ $2.6$ $0.38\pm 0.58$ $2.5$ $0.23\pm 0.59$ $2.5$
$C_{qq}^{(3)[3333]}$ $0.040\pm 0.020$ $10.4$ $1.0\pm 1.4$ $1.5$ $1.4\pm 1.2$ $1.5$ $0.6\pm 1.4$ $1.5$
@@
$C_{qq}^{(3)[aabb]}$ $0.032\pm 0.019$ $11.2$ $0.032\pm 0.078$ $6.9$ $0.031\pm 0.079$ $7.0$ $0.031\pm 0.079$ $7.0$
$C_{qq}^{(3)[abba]}$ $0.036\pm 0.018$ $11.1$ $0.06\pm 0.20$ $4.5$ $0.06\pm 0.20$ $4.4$ $0.06\pm 0.20$ $4.4$
$C_{lq}^{(1)[aa33]}$ $0.019\pm 0.035$ $10.1$ $0.28\pm 0.32$ $3.1$ $0.35\pm 0.30$ $3.2$ $0.33\pm 0.33$ $3.1$
$C_{lq}^{(1)[aabb]}$ $0.024\pm 0.030$ $10.4$ $0.035\pm 0.056$ $8.0$ $0.028\pm 0.055$ $8.0$ $0.030\pm 0.056$ $8.1$
$C_{lq}^{(3)[aa33]}$ $0.022\pm 0.034$ $10.0$ $0.00\pm 0.25$ $4.2$ $0.02\pm 0.22$ $4.7$ $0.01\pm 0.25$ $4.2$
$C_{qe}^{[33aa]}$ $0.020\pm 0.035$ $9.9$ $0.58\pm 0.45$ $2.5$ $0.59\pm 0.40$ $2.6$ $0.55\pm 0.46$ $2.5$
$C_{qe}^{[aabb]}$ $0.015\pm 0.032$ $10.8$ $0.055\pm 0.099$ $6.0$ $\pm 0.10 6.06.0 0.0$$05\pm 0.10$ $6.0$
$C_{qu}^{(1)[3333]}$ $0.08\pm 0.36$ $3.4$ $0.7\pm 1.3$ $1.7$ $0.6\pm 1.2$ $1.7$ $0.7\pm 1.3$ $1.7$
$C_{qu}^{(1)[aa33]}$ $0.02\pm 0.34$ $3.6$ $1.0\pm 2.2$ $1.3$ $1.5\pm 2.3$ $1.2$ $1.1\pm 0.0$ $1.3$
@@ 10 TeV Individual 10 TeV Global
$C_{i}$ $\Lambda/\sqrt{|C_{i}|}$ $C_{i}$ $\Lambda/\sqrt{|C_{i}|}$
@@
$C_{\phi u}^{[aa]}$ $4.8\pm 4.3$ $2.8$ $\color[rgb]{1,0,0}-8\pm 7$}
$C_{\phi d}^{[33]}$ $\color[rgb]{1,0,0}-10.9\pm 4.2$} $\color[rgb]{1,0,0}-8.1\pm 6.7$}
@@
$C_{ll}^{[aabb]}$ $2.5\pm 2.6$ $3.7$
$C_{ll}^{[abba]}$ $0.1\pm 1.0$ $6.8$ $0.6\pm 3.3$ $3.8$
$C_{qq}^{(1)[3333]}$ $0.45\pm 0.23$ $10.7$
$C_{qq}^{(1)[aa33]}$ $0.235\pm 0.099$ $15.1$
$C_{qq}^{(1)[a33a]}$ $0.214\pm 0.096$ $15.5$
$C_{qq}^{(1)[aabb]}$ $0.43\pm 0.22$ $10.2$
$C_{qq}^{(1)[abba]}$ $0.47\pm 0.22$ $10.7$
$C_{qq}^{(3)[3333]}$ $0.56\pm 0.23$ $10.2$
$C_{qq}^{(3)[aa33]}$ $0.22\pm 0.13$ $13.5$ $0.5\pm 3.2$ $4.2$
$C_{qq}^{(3)[a33a]}$ $0.18\pm 0.11$ $15.7$ $3.8\pm 4.2$ $2.8$
$C_{qq}^{(3)[aabb]}$ $0.36\pm 0.21$ $11.5$
$C_{qq}^{(3)[abba]}$ $0.40\pm 0.21$ $10.6$
@@
$C_{lq}^{(1)[aa33]}$ $0.21\pm 0.38$ $10.1$ $\color[rgb]{1,0,0}-7.0\pm 8.1$}
$C_{lq}^{(1)[33aa]}$ $3.6\pm 2.5$ $3.4$ $2.6\pm 8.2$
$C_{lq}^{(1)[aabb]}$ $0.17\pm 0.31$ $11.3$ $3.9\pm 4.6$ $2.8$
$C_{lq}^{(3)[3333]}$ $8.2\pm 5.8$
$C_{lq}^{(3)[aa33]}$ $0.25\pm 0.39$ $9.7$ $5.6\pm 7.0$
$C_{lq}^{(3)[33aa]}$ $0.21\pm 0.29$ $11.1$ $0.2\pm 1.2$ $7.3$
$C_{lq}^{(3)[aabb]}$ $0.011\pm 0.056$ $29.2$ $0.1\pm 1.1$ $7.7$
$C_{ee}^{[aa33]}$ $0.0\pm 2.1$ $4.9$ $2.0\pm 5.2$ $2.8$
$C_{ee}^{[aabb]}$ $1.4\pm 1.5$ $4.8$ $2.4\pm 5.1$ $3.2$
@@
$C_{le}^{[aa33]}$ $4.2\pm 7.1$
$C_{le}^{[33aa]}$ $\color[rgb]{1,0,0}-6.7\pm 7.7$}
$C_{le}^{[aabb]}$ $0.4\pm 4.2$ $3.3$ $2.1\pm 4.8$ $3.2$
@@
$C_{qe}^{[33aa]}$ $0.22\pm 0.38$ $10.0$
$C_{qe}^{[aabb]}$ $0.23\pm 0.36$ $10.4$
$C_{qu}^{(1)[3333]}$ $0.3\pm 3.1$ $3.9$ $\color[rgb]{1,0,0}7.5\pm 7.5$}
$C_{qu}^{(1)[aa33]}$ $0.1\pm 2.9$ $4.2$
$C_{qu}^{(1)[33aa]}$ $\color[rgb]{1,0,0}-7.0\pm 6.9$}
$C_{qu}^{(1)[aabb]}$ $3.6\pm 9.1$
@@
$C_{quqd}^{(1)[3333]}$ $0.10\pm 0.40$ $10.2$
$C_{quqd}^{(8)[3333]}$ $0.16\pm 0.48$ $9.2$
$C_{lequ}^{(1)[3333]}$ $1.2\pm 3.7$ $3.5$